• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 11:03:08 on January21,2008

    Entry number 226565

    keyword=A2 HV being changed

    I just got the display for A2 ADC right yesterday so they were left at low HV. I will change the HV for A2 to get it to a reasonable level to be useful for the production data. A few run where we have many pions would be good to check the calibration of A2 before production. I will make another HALOG entry when HV will be stable.